# Direct measurement of isotope shifts in the barium 6s2 1S0 - 5d6p 3D transition

### v2

@@ Appendix C The linewidths of the fluorescence spectrum
To estimate the effect of transverse velocity distribution on the spectrum linewidth, we first measured the velocity distribution of the atomic beam. We measured the atomic fluorescence spectrum at a point 2.5mm shifted from the normal incident point along the excitation laser direction, where the incident angle between the excitation laser and atoms was 0.35 rad. By analyzing the Doppler shifts of the spectrum and given geometry between the laser beam and atomic beam, the fit of spectral data to the Boltzmann distribution yielded the root-mean-squared atomic velocity of 4352 m/s and temperature of 10508 K (Fig. 7). From the derived velocity distribution, the transverse velocity profile of a collimated effusive atomic beam could be derived from the model below [30]
@@
where, is an incomplete gamma function, is the oven aperture size, is the detection area size, , and with the distance between the oven and detection area. Figure 8 shows the transverse velocity distribution derived from the velocity distribution, setup geometry, and corresponding Doppler shifts. The velocity distribution’s full width at half maximum (FWHM) is 6.150.03 m/s, equivalent to Doppler shifts of 14.80.1 MHz.
@@
The spectra with different excitation laser intensities were obtained to confirm the absence of power broadening, and no increase in spectral width was observed (Fig. 9). In addition, the interaction time between the laser field and atoms was about 3 s, and the corresponding transit time broadening was negligible.
Appendix D Comparison of different peak models
@@
Table 4 shows the fit results with various fit models. Although the detailed peak shape differs by the models, the extracted isotope shift values agree within 1 MHz and the choice of peak model does not affect the main result of this work. With pure Voigt peak model (), the fitted peak is mostly Lorentzian, while the uncertainty of Gaussian width is large. We could fit the data well with Voigt with background Gaussian model () while the Lorentizan width is fixed to the natural linewidth ( MHz), although the mean squared error (MSE) was slightly higher than the unfixed case. The fitted Gaussian FWHM width is MHz, comparable with expected Doppler broadening calculated in Appendix C
| Fit parameters | ( fixed) | ||
| (MHz) | 391.3 1.3 | 392.5 0.6 | 392.5 0.8 |
| (MHz) | 176.7 0.9 | 177.8 0.4 | 178.2 0.5 |
| (MHz) | 399.2 2.3 | 400.6 1.0 | 401.2 1.3 |
| (MHz) | 118.6 3.0 | 125.9 1.4 | 126.6 1.4 |
| (MHz) | 14.2 0.3 | 11.2 0.4 | 4.58 |
| (MHz) | 0.1 21.0 | 3.1 0.5 | 9.0 0.2 |
| (MHz) | N/A | 69.8 2.4 | 60.4 1.3 |
| N/A | 0.33 0.02 | 0.68 0.02 | |
| MSE | 1.03 | 2.31 | 6.16 |

### v3

@@ Appendix C Relative uncertainty of the wavelength meter measurement
To estimate the effect of transverse velocity distribution on the spectrum linewidth, we first measure the velocity distribution of the atomic beam. We measure the atomic fluorescence spectrum at a point 2.5mm shifted from the normal incident point along the excitation laser direction, where the incident angle between the excitation laser and atoms is 0.35 rad. By analyzing the Doppler shifts of the spectrum and given geometry between the laser beam and atomic beam, the fit of spectral data to the Boltzmann distribution yielded the root-mean-squared atomic velocity of 4352 m/s and temperature of 10508 K (Fig. 7). From the derived velocity distribution, the transverse velocity profile of a collimated effusive atomic beam could be derived from the model below [35]
@@
where, is an incomplete gamma function, is the oven aperture size, is the detection area size, = kT/m, and with the distance between the oven and detection area. Figure 8 shows the transverse velocity distribution derived from the velocity distribution, setup geometry, and corresponding Doppler shifts. The velocity distribution’s full width at half maximum (FWHM) is 6.150.03 m/s, equivalent to Doppler shifts of 14.80.1 MHz.
@@
The spectra with different excitation laser intensities are obtained to confirm the absence of power broadening, and no increase in spectral width is observed (Fig. 9). In addition, the interaction time between the laser field and atoms was about 3 s, and the corresponding transit time broadening is negligible.
Appendix E Comparison of different peak models
@@
Table 5 shows the fit results with various fit function models. Among the various models, Lorentzian peak with Gaussian background () shows the lowest reduced of 1.56, suggesting the model agrees the most with the data. Although the detailed peak shape differs by the models, the extracted isotope shift values agree within 1 MHz and the choice of peak model does not affect the main result of this work. With the pure Voigt peak model (), the fitted peak is mostly Lorentzian, while the uncertainty of the Gaussian width is large. We could also fit the data with the Voigt with the background Gaussian model () while the Lorentizan width is fixed to the natural linewidth ( MHz), although the mean squared error (MSE) is slightly higher than the unfixed case. The fitted Gaussian FWHM width is MHz, comparable with the expected Doppler broadening calculated in Appendix D.
| Fit parameters | ( fixed) | |||
| (MHz) | 392.9 0.6 | 391.9 1.2 | 392.9 0.6 | 392.7 0.8 |
| (MHz) | 178.1 0.4 | 176.3 0.9 | 178.1 0.4 | 178.3 0.5 |
| (MHz) | 401.4 1.0 | 401.3 2.2 | 401.4 1.0 | 401.4 1.4 |
| (MHz) | 124.3 1.2 | 122.3 2.3 | 125.3 1.2 | 126.7 1.4 |
| (MHz) | 12.3 0.2 | 15.5 0.5 | 12.3 0.6 | 4.58 (fixed) |
| (MHz) | N/A | 0.1 42.6 | 0.1 29.2 | 9.0 0.2 |
| (MHz) | 73.3 2.5 | N/A | 73.3 2.7 | 60.3 1.3 |
| 0.30 0.02 | N/A | 0.30 0.03 | 0.68 0.02 | |
| MSE | 3.61 | 1.35 | 3.6 | 6.06 |
| Reduced | 1.56 | 4.69 | 1.57 | 2.45 |